OpenStudy (anonymous):

How do I answer this question 3(x+1)-8=14-2(3x-4)

OpenStudy (texaschic101):

your solving for x 3(x + 1) - 8 = 14 - 2(3x - 4) --- distribute through the parenthesis 3x + 3 - 8 = 14 - 6x + 8 --- combine like terms 3x - 5 = -6x + 22 --- add 6x to both sides 3x + 6x - 5 = 22 --- add 5 to both sides 3x + 6x = 22 + 5 --- combine like terms 9x = 27 -- divide both sides by 9 x = 3